OpenStudy (anonymous):

Simplify (16x^6y)^(1/2)

OpenStudy (anonymous):

It is written in the form of \[(16x^{6}y)^{1/2}\]

OpenStudy (lgbasallote):

\[\LARGE (a^b)^c = a^{bc}\]

OpenStudy (lgbasallote):

\[\LARGE a^b c)^d = a^{bd}c^d\]

OpenStudy (anonymous):

Is it \[4\sqrt{x^3y}\]

OpenStudy (lgbasallote):

lol did you get what those meant :p

OpenStudy (lgbasallote):

not exactly... \[\LARGE (16x^6y)^{1/2} = (16^{1/2}) \times ( x^{6/2}) \times (y^{1/2})\]

OpenStudy (lgbasallote):

you're almost right just that x^3 shouldnt be in the root

OpenStudy (anonymous):

Oh I see it! =) Thank you so much! =)